NATURE OF TORSION-INVERSION COUPLING IN CH_{3}NH_{2}, CH_{3}OH_{2}^{+} AND CH_{3}CH_{2}·

\underline{RAM S BHATTA}, AMY GAO AND DAVID S PERRY, Department of Chemistry, The University of Akron, OH 44325-3601.

Two-dimensional torsion-inversion surfaces for methylamine, protonated methanol and ethyl radical were calculated and fit to a function containing a polynomial in the inversion angle(τ) and trigonometric functions of the torsional angle(α). Calculations were done at the B3LYP, MP2, and CCSD(T) levels with the 6-311++G(d,p) and 6-311++G(3df, 2p) basis sets and partial optimization. CH_{3}NH_{2}, CH_{3}OH_{2}^{+} and CH_{3}CH_{2}· have G_{12} symmetry with 6-equivalent minima which are located by the various calculations at inversion angles 6.5 to 11; 42 to 45.5 and 52.5 to 55 degrees respectively on either side of planar. The three molecules have very different barriers to inversion ranging from no barrier for CH_{3}CH_{2}· to 838 cm^{−1} for CH_{3}OH_{2}^{+} to 1837 cm^{−1} for CH_{3}NH_{2}. The dominant torsion-inversion coupling term in all cases has the form τcos3α.